OpenStudy (anonymous):

what is the integral e^x/1+e^2x dx

OpenStudy (anonymous):

\[\int\limits_{?}^{?} e^x/e^2x\]

OpenStudy (lgbasallote):

i have a feeling this has trigonometric substitution yes?

OpenStudy (anonymous):

i think so but i dont know how to do it, sorry its \[\int\limits_{?}^{?} e^{x}/1+e^{2x}\]

OpenStudy (lgbasallote):

hmmm i think i can solve the first few parts....

OpenStudy (anonymous):

its a multiple choice, so there are possible solution presented should I type them for you?

OpenStudy (lgbasallote):

let e^x = u u^2 = e^2x du = e^xdx du/(1+u^2)  tan^2 (theta) = 1 + u^2 sin (theta) = u du = cos (theta) d(theta) cos (theta)/tan^2 (theta) hmm..any ideas?

OpenStudy (anonymous):

\[ \arctan (e^x)\]

OpenStudy (anonymous):

The derivative of \[ \arctan (u)\] with respect ot \[ u\] is \[ \frac 1 {1+u^2}\]

OpenStudy (anonymous):

the antiderivative is \[e^x+(1/2)e^{2x}+c\] that is if you mean \[\int\limits_{}^{}(e^x/1)+e^{2x}dx\]

OpenStudy (anonymous):

@ anonymoustwo: no thats not what I mean but eliassaab answered me thank you. I didnt know that the derivatite of 1/1_u^2 is arctan u thank you guys for the help all of you